OpenStudy (anonymous):

Divide. Reduce your answer to lowest terms. 14/15 divided By 7/9

OpenStudy (anonymous):

@Panda.exe

OpenStudy (panda.exe):

1 1/5

OpenStudy (anonymous):

i didnt write a sentence.. i just did Is it still 1 1/5?

OpenStudy (panda.exe):

what do u mean

OpenStudy (anonymous):

i didnt write a part of the problem

OpenStudy (anonymous):

Which was: R

OpenStudy (anonymous):

reduce your answer to lowest terms

OpenStudy (panda.exe):

14/15 / 7/9 = (14×9) / (15×7) = 126/105 = 6/5 = 1.2

OpenStudy (panda.exe):

1/2

OpenStudy (anonymous):

but i dont have that as an answer choice I forgot to write them again :/

OpenStudy (panda.exe):

face palm

OpenStudy (anonymous):

A. 5/6 B. 1 1/5 C. 2/3 D. 2 3/5

OpenStudy (panda.exe):

b

OpenStudy (anonymous):

ok thx

OpenStudy (panda.exe):

fan me bro

OpenStudy (anonymous):

i am a girl :'(

OpenStudy (panda.exe):

brolet xD

OpenStudy (anonymous):

But ok becuase you are being a great help! :D

OpenStudy (anonymous):

lol

OpenStudy (panda.exe):

and so u any time u need help tag me ok